Larry can paint a room in 8 hours by himself. After he worked for 2 hours, Patrick joins him and they finish the job in 2 hours. How long would it take Patrick to do the entire job without Larry?: Larry can paint a room in 8 hours by himself. After he worked for 2 hours, Patrick joins him and they finish the job in 2 hours. How long would it take Patrick to do the entire job without Larry?
Answer by stanbon(19722)

Larry can paint a room in 8 hours by himself. After he worked for 2 hours, Patrick joins him and they finish the job in 2 hours. How long would it take Patrick to do the entire job without Larry?
------------------------------
Larry DATA:
time = 8 hr/job ; rate = (1/8)job/hr.
In 2 hrs Larry completes 2(1/8) = 1/4th of the job.
-------------------------------------------
Patrick DATA:
time = x hr/job ; rate = 1/x job/hr.
EQUATION:
time*rate = work done
2*((1/8)+(1/x)) = 3/4
(1/8)+(1/x)= 3/8
1/x = 2/8
x = 4 hrs (time for Patrick to do the work alone)
=================================================
Cheers,
Stan H.